Suppose a firm has two plants. The costs are:

TC1 = 40 + 12Q1 + Q12

TC2 = 30 + 10Q2 + .5Q22

The firm’s demand is Pm = 60 - .6Qm.

Find the firm’s

  1. profit-maximizing P* and Q*,
  2. outputs allocated to plants 1 and 2,
  3. overall profit.
Answer #1

Lets do this question by (ouonot Couonot duopoly Foo firmi P & 6o - 0.6om 1 Om = Qet az BE bo - 0.0Q,-0.6.Q2TR-TC profit (IT) Pmx (9,792) - (Tat Ter) = (x1.41 x 31)-(303-637 4 20:45) = u 고 1 283.71 - 724.08 A = $500 (approxy